# Learning High-Order Relationships of Brain Regions

### 02 Dec 2023 v1

@@ 6 Conclusion
In this work, we propose HyBRiD for identifying maximally informative yet minimally redundant (MIMR) high-order relationships of brain regions. To effectively optimize our model, we further proposed a novel information bottleneck framework with theoretical guarantees. Our method outperforms state-of-the-art models of hypergraph construction and connectivity-based prediction. The result analysis shows the effectiveness of our model. We expect such advancements could benefit clinical studies, providing insights into neurological disorders, and offering improved diagnostic tools in neurology.
@@
HyBRiD only considers static high-order relations. Given that ABCD tasks are dynamic, including temporal changes and interactions, it will be interesting to study the evolution of these high-order relationships.
@@ Appendix A Proof of the Upper Bound
According to Lemma 1, we have $I(X;Y|Z)\leq I(X;Y)$, which finishes the proof.
Theorem 2.
@@
Figure 7: The graphical model of random variables $\bm{h}^{k}$ and $X$
@@
According to the definitions of $X$ and $H$, which are described in Section 4, we can draw a graphical model of them in Figure 7. Define a new random variable $\bm{h}^{k_{1}:k_{2}}=[\bm{h}^{k_{1}},\bm{h}^{k_{1}+1},\cdots,\bm{h}^{k_{2}-1},\bm{h}^{k_{2}}]$, which is a concatenation from $\bm{h}^{k_{1}}$ to $\bm{h}^{k_{2}}$. According to Corollary 1.1 we have
$I(H;X)$ $\leq I(\bm{h}^{1},X)+I(\bm{h}^{2:K},X)$ (19)
$\leq I(\bm{h}^{1},X)+I(\bm{h}^{2},X)+I(\bm{h}^{3:K},X)$
$\leq I(\bm{h}^{1},X)+I(\bm{h}^{2},X)+I(\bm{h}^{3},X)+\cdots$
$\leq\sum_{k=1}^{K}I(\bm{h}^{k};X)$
Theorem 3.
@@
$I(\bm{h}^{k};X)\leq\sum_{i=1}^{N}I(\bm{h}^{k}_{i};X_{i})=\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i}),\quad\forall 1\leq k\leq K$ (21)
And this is exactly the conclusion in Kim et al. (2021) if we consider $\bm{h}^{k}$ and $X$ as $X$ and $Z$ respectively in their paper. ∎
@@ Appendix B Connectome-Based Predictive Modeling
Shen et al. (2017); Finn et al. (2015)) has shown tremendous promise in recent years in detecting imaging biomarkers. (Rosenberg et al., 2015; Dubois et al., 2018a; Rosenberg et al., 2020; 2016). Such a model, based chiefly on functional MRI data, can measure the significance of the input edge weights, which is revealed by a correlation coefficient that reflects the correlation between the edge weights and the neurological outcomes. One could expect a large correlation coefficient to indicate the high quality of edge weights. We utilize the CPM as an evaluation model to evaluate the quality of our learned hyperedges. Here is a pipeline overview of the CPM process:
@@
Since positive edges and negative edges will cancel out with each other when being summed, we adopt the combining strategy in Boyle et al. (2023).
@@
According to step 6, the evaluation of the predictive model could be measured by the correlation between predicted and ground-truth outcomes Shen et al. (2017). Since CPM is a linear model that predicts the outcome based on the sum of significant edge weights, the correlation is equal to the correlation between the sum and ground-truth outcomes (which is exactly the r in Eq. 13). Hence, one can expect a larger correlation if the edge weights are more correlated (and thus are more predictive).
@@
Appendix C Dataset details
@@
Dataset Rest 1 SST 1 EN-back 1 MID 1 Rest 2 SST 2 EN-back 2 MID 2
#(instances) 1676 1673 1678 1678 1949 1053 1044 1062
length of time series 375 437 362 403 375 437 362 403
Table 3: Statistics of 8 datasets.
Appendix D Training Details
@@
We train our model on a machine with a Intel Xeon Gold 6326 CPU and RTX A5000 GPUs.
@@
See Table 4 for software we used and the versions.
@@
Table 4: Software versions
@@
The hyperparameters selection is shown in Table 5. Some crucial hyperparameters ablation experiments can be found in Appendix E.
notation meaning value
$lr$ learning rate $1\times 10^{-3}$
$K$ number of hyperedges $32$
$\beta$ trade-off coefficients information bottleneck $0.2$
$[h_{1},h_{2},h_{3}]$ hidden sizes of the dim reduction MLP $[32,8,1]$
$B$ batch size $64$
Table 5: Hyperparameter choices
Appendix E More Ablation Studies
E.1 Parameters Sharing
As previously discussed, the scarcity of fMRI data can result in serious over-fitting. In our experiments, we observed that when the model is trained on each dataset individually, the validation set loss remains stagnant, which indicates over-fitting. To address this, we examined the model’s performance under various parameter-sharing strategies. Specifically, we have three learnable components: the shallow embeddings in Equation 3, the dim reduction MLP in 6 and the LinearHead $\mathcal{F}_{l}$. These components can be either shared or condition-specific. Results of different parameter-sharing strategy is shown in Table 6.
Strategy ID ShallowEmbeddings DimReduction LinearHead Average Score
1 $0.337$
2 $0.335$
3 $0.508$
4 $\underline{0.539}$
5 $0.351$
6 $0.384$
7 $0.506$
8 $\mathbf{0.558}$
Table 6: The table summarizes the model’s average performance under various parameter-sharing strategies. The symbol ✓ indicates a component being condition-specific, while an empty cell implies a component being shared across conditions.
Note that strategy 1 is equivalent to training models separately for each condition. From the table, we find that the DimReduction component plays the most significant role. Making DimReduction condition-specific enhances the average performance by $10\%-20\%$. Strategy 8 is the best. However, domain experts may expect different structures under different conditions. Therefore, we still adopt condition-specific ShallowEmbeddings (i.e. strategy 4), given that their scores are really close. We argue that the fact that strategy 4 is not better than strategy 8 is because our prediction target is fluid intelligence, which is a general measurement of cognition and thus tends to be condition-agnostic. We anticipate a potential shift in results if we switch to a more condition-specific cognition phenotype outcome.
E.2 Choices of number of hyperedges $K$
@@
E.3 Choices of the trade-off coefficient $\beta$
@@
Appendix F Runtime Comparison
@@
Appendix G More Visualizations
@@
We visualize the most significant hyperedge under the EN-back condition. We observe a coordinated interaction of numerous brain regions, each fulfilling specific roles. Notably, some of these regions serve multi-functional purposes:
@@
  • Visual Processing: Calcarine_R, Lingual_L, Fusiform_L. These regions are responsible for visual perception and some of them are related to complex visual contents like symbols and human faces, which were presented during the task.
@@
Figure 9: Visualization of the most significant hyperedge in EN-back condition
Resting-state brain region importance
Figure 10: Region importance of resting state.
body>

### 13 Feb 2024 v2

@@ 6 Conclusion
In this work, we proposed HyBRiD for identifying maximally informative yet minimally redundant (MIMR) high-order relationships of brain regions. To effectively optimize our model, we further proposed a novel information bottleneck framework and derived its theory. Our method outperforms state-of-the-art models. The result analysis shows the effectiveness of our model. We expect such advancements could benefit clinical studies, providing insights into neurological disorders, and offering improved diagnostic tools in neurology and other related fields.
@@
HyBRiD only considers static high-order relations. Given that fMRI tasks are dynamic, including temporal changes and interactions, it will be interesting to study the evolution of these high-order relationships.
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Proof of the Upper Bound
According to Lemma A.1, we have $I(X;Y|Z)\leq I(X;Y)$, which finishes the proof.
Theorem A.3.
@@
Figure 7: The graphical model of random variables $\boldsymbol{h}^{k}$ and $X$
@@
According to the definitions of $X$ and $H$, which are described in Section 4, we can draw a graphical model of them in Figure 7. Define a new random variable $\boldsymbol{h}^{k_{1}:k_{2}}=[\boldsymbol{h}^{k_{1}},\boldsymbol{h}^{k_{1}+1},\cdots,\boldsymbol{h}^{k_{2}-1},\boldsymbol{h}^{k_{2}}]$, which is a concatenation from $\boldsymbol{h}^{k_{1}}$ to $\boldsymbol{h}^{k_{2}}$. According to Corollary A.2 we have
$I(H;X)$ $\leq I(\boldsymbol{h}^{1},X)+I(\boldsymbol{h}^{2:K},X)$ (19)
$\leq I(\boldsymbol{h}^{1},X)+I(\boldsymbol{h}^{2},X)+I(\boldsymbol{h}^{3:K},X)$
$\leq I(\boldsymbol{h}^{1},X)+I(\boldsymbol{h}^{2},X)+I(\boldsymbol{h}^{3},X)+\cdots$
$\leq\sum_{k=1}^{K}I(\boldsymbol{h}^{k};X)$
Theorem A.4.
@@
$I(\boldsymbol{h}^{k};X)\leq\sum_{i=1}^{N}I(\boldsymbol{h}^{k}_{i};X_{i})=\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i}),\quad\forall 1\leq k\leq K$ (21)
And this is exactly the conclusion in (Kim et al., 2021) if we consider $\boldsymbol{h}^{k}$ and $X$ as $X$ and $Z$ respectively in their paper. ∎
@@ Appendix B Connectome-Based Predictive Modeling
Shen et al. (2017); Finn et al. (2015) have shown tremendous promise in recent years in detecting imaging biomarkers by CPM (connectome-based predictive modeling) (Rosenberg et al., 2015; Dubois et al., 2018; Rosenberg et al., 2020, 2016). Such a model, based chiefly on functional MRI data, can measure the significance of the input edge weights, which is revealed by a correlation coefficient that reflects the correlation between the edge weights and the neurological outcomes. One could expect a large correlation coefficient to indicate the high quality of edge weights. We utilize the CPM as an evaluation model to evaluate the quality of our learned hyperedges. Here is a pipeline overview of the CPM process:
@@
Since positive edges and negative edges will cancel out with each other when being summed, we adopt the combining strategy in (Boyle et al., 2023).
@@
According to step 6, the evaluation of the predictive model could be measured by the correlation between predicted and ground-truth outcomes (Shen et al., 2017). Since CPM is a linear model that predicts the outcome based on the sum of significant edge weights, the correlation is equal to the correlation between the sum and ground-truth outcomes (which is exactly the $r$ in Equation 13). Hence, one can expect a larger correlation if the edge weights are more correlated (and thus are more predictive).
@@
Appendix C Computational Complexity
Suppose we have $N$ regions as $N$ nodes and $K$ hyperedges. For the Constructor, the unignorable computation is from the mask operation. The computational complexity of this step for each hyperedge is $O(Nd)$ since it does a pairwise multiplication operation of a matrix of size $N\times d$. Given that there are $K$ hyperedges, the total complexity is $O(NKd)$. For the Weighter, the computation is from the dim reduction operation and the linear head. The dim reduction operation is an MLP. In this work, the hidden dimensions are a fraction of the original feature dimension. Therefore, the complexity of the dim-reduction MLP is $O(d^{2})$. The linear head only contributes $O(d)$, which is neglectable. As a result, the computational complexity of the whole model is $O(NKd+d^{2})=O(N^{2}K)$ since the feature dimension is equal to the number of regions (See Appendix D for features we used). This complexity is just at the same scale as that of MLPs even though we are addressing a more challenging task: identifying high-order relationships in an exponential space.
Appendix D Dataset details
D.1 Preprocessing of ABCD Dataset
We use the preprocessed ABIDE dataset from the official website. We preprocessed the restricted ABCD dataset ourselves. Below is the preprocessing procedure of the ABCD dataset.
raw data to voxel-level fMRI time series
The fMRI data is processed using BioImage Suite (Joshi et al., 2011). First, we performed motion correction and slice-time correction using SPM5; and via BioImage Suite, the data were registered to a standardized $3mm\times 3mm\times 3mm$ common space, where we generated masks representing white matter, gray matter, and cerebrospinal fluid (CSF) and computed the mean time courses for both white matter and CSF. We orthogonalized each gray matter time course with respect to the mean time courses of both white matter and CSF, and we orthogonalized each gray matter time course to the six motion-related signals via SPM5. We then applied a bandpass Butterworth filter with a frequency range of $0.02$Hz to $0.1$Hz to the orthogonalized time courses. We used a Gaussian kernel with a full-width at half-maximum (FWHM) of $6$mm to enhance spatial coherence and spatial smoothing. Lastly, we removed the linear trend from all signals in accordance with the methodology detailed in (Shen et al., 2013). We deleted scans with more than 0.10 mm mean frame-to-frame displacement. Additional details about the standard preprocessing procedures, such as slice time and motion correction, and registration to the MNI template can be found in (Greene et al., 2018) and (Horien et al., 2019).
voxel-level fMRI time series to region features
The fMRI time series data of a human subject is represented in four dimensions (3 spatial dimensions + 1 temporal dimension), which can be imagined as a temporal sequence of 3D images. First, brain images are parceled into regions (or nodes) using the AAL3v1 atlas (Rolls et al., 2020). Following previous works (Kan et al., 2022b; Li et al., 2021; Thomas et al., 2022), each region’s time series is obtained by averaging all voxels in that region. Consistent with previous connectivity-based methods (Li et al., 2021; Kan et al., 2022b; Ktena et al., 2018; Said et al., 2023), for each region, we use its Pearson correlation coefficients to all regions as its features. We randomly split the data into train, validation, and test sets in a stratified fashion. The split ratio is 8:1:1.
D.2 Dataset Statistics
@@
Dataset ABIDE ABCD
FIQ VIQ PIQ Rest 1 SST 1 EN-back 1 MID 1 Rest 2 SST 2 EN-back 2 MID 2
#(instances) 1035 1035 1035 1676 1673 1678 1678 1949 1053 1044 1062
length of time series 196 196 196 375 437 362 403 375 437 362 403
Table 4: Statistics of datasets we use. ABIDE dataset only contains resting-state fMRI data. ABCD contains resting-state and task-based data. We use ABCD of $2$ timepoints. For example, Rest 1 means resting-state fMRI data from timepoint $1$.
Appendix E Training Details
Due to the data scarcity, we found training on single sub-datasets of the ABCD dataset leads to severe overfitting. To mitigate this, we train our model as well as all the baselines on all datasets together and report the results individually. Note that on the ABIDE dataset, we train our model under the three targets separately since we don’t encounter such an issue on the ABIDE dataset.
@@ Appendix E Training Details
We train our model on a machine with an Intel Xeon Gold 6326 CPU and RTX A5000 GPUs.
@@
See Table 5 for the software we used and the versions.
@@
Table 5: Software versions.
@@
The hyperparameters selection is shown in Table 6. Some crucial hyperparameters ablation experiments can be found in Appendix H.
notation
meaning
value
$lr$
learning rate
$1\times 10^{-3}$
$K$
number of hyperedges
$32$
$\beta$
trade-off coefficients information bottleneck
$0.2$
$[h_{1},h_{2},h_{3}]$
hidden sizes of the dim reduction MLP
$[32,8,1]$
$B$
batch size
$64$
Table 6: Hyperparameter choices.
Appendix F Why is HyBRiD not the Best on the Rest 1 Dataset?
Our model reaches the second-best on the Rest 1 dataset. To find out the reason, we inspect the performance on both the training, validation and test datasets. We summarize the performances of our model and the SOTA one in the table below
model train validation test
BrainNetTF $0.967$ $0.484$ $0.334$
HyBRiD (Ours) $0.984$ $0.491$ $0.223$
Table 7: Our model’s performance compared to the SOTA one on the Rest 1 dataset.
We find our model outperforms BrainNetTF on both training and validation datasets but fails on test datasets. Therefore, given that Rest 1 is the most noisy dataset, we argue that the imperfection is because
1) Our model overfits the training dataset more than BrainNetTF.
2) The validation score is higher, indicating the reason is the discrepancy between the training dataset and the test dataset.
Appendix G Evaluation on Synthetic Dataset
Dataset Synthesis
Metric $K=1$ $K=5$ $K=10$ $K=30$
P R F1 P R F1 P R F1 P R F1
Random $0.158$ $0.094$ $0.110$ $0.182$ $0.170$ $0.171$ $0.224$ $0.218$ $0.218$ $\underline{0.288}$ $0.250$ $0.266$
$k$NN $0.533$ $\mathbf{1.000}$ $0.696$ $0.247$ $0.241$ $0.243$ $0.144$ $0.156$ $0.148$ $0.139$ $0.128$ $0.133$
$l_{1}$ hypergraph $0.160$ $\mathbf{1.000}$ $0.276$ $0.180$ $0.016$ $0.028$ $0.135$ $0.016$ $0.028$ $0.256$ $0.026$ $0.047$
$l_{2}$ hypergraph $\underline{0.987}$ $0.825$ $\underline{0.897}$ $\underline{0.412}$ $\underline{0.631}$ $\underline{0.494}$ $\underline{0.276}$ $\underline{0.499}$ $\underline{0.351}$ $0.233$ $\underline{0.470}$ $0.310$
HGCN $0.625$ $\mathbf{1.000}$ $0.769$ $0.196$ $0.477$ $0.278$ $0.188$ $0.379$ $0.251$ $0.277$ $0.364$ $\underline{0.314}$
HyBRiD (Ours) $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{0.595}$ $\mathbf{0.680}$ $\mathbf{0.631}$ $\mathbf{0.504}$ $\mathbf{0.604}$ $\mathbf{0.549}$ $\mathbf{0.428}$ $\mathbf{0.393}$ $\mathbf{0.401}$
Table 8: Precision, recall and F1 score on synthetic dataset under different $K$, where $K$ is the number of hyperedges. The performance of the Random method increases as $K$ increases because the Hungarian algorithm is likely to provide better matches when there are more candidates.
Our synthetic dataset is constructed as follows:
1) Structure Generation: For each hyperedge, we randomly sample the hyperedge degree $d$ from a discrete uniform distribution between $2$ and $d_{\text{max}}$. After that, we randomly sample $d$ nodes as members of this hyperedge.
2) Feature generation: For each hyperedge, we randomly sample a scalar value $v$ from the uniform distribution $U(0,1)$. We then randomly sample the features from $U(0,2v)$ for each node in this hyperedge.
3) Label Generation: For each hyperedge, we calculate the maximum of its node features as the summary of the hyperedge. We sum the summaries of all hyperedges to get a single value $Y$ as the label of the hypergraph.
Note that consistent with our settings of real-world datasets, the structure is shared across all hypergraphs, while node features and hyperedge weights are different.
Metrics
We use the macro-precision, recall and F1 score to measure the correctness of the learned hyperedges with respect to the ground-truth ones. Note that when learning on the synthetic dataset, the order of hyperedge may differ from the ground truths. Therefore, the Hungarian algorithm is employed to match the learned hyperedges with the ground truth. The precision, recall and F1 score are calculated after matching.
Baselines
We use the hyperedge construction methods in Section 5, i.e. $k$NN(Huang et al., 2009), $l_{1}$ hypergraph(Wang et al., 2015) and $l_{2}$ hypergraph(Jin et al., 2019) as our baselines. Besides, we implement a hypergraph structure learning model. This is a hypergraph convolutional neural network (HGCN) (Bai et al., 2021), where both the parameters of layers and the hypergraph structure (the incidence matrix) are learnable.
Implementaton Details
The maximum degree $d_{\text{max}}$ is set to $34$, which is the maximum degree of hyperedges of the ABCD dataset according to the analysis in Section 5. The number of nodes is set to $164$, which is the number of regions of AAL3v1 atlas, used in the ABCD experiment. We conduct the experiments under different numbers of hyperedges (i.e. $K=1,5,10,30$).
Results
We report the results under different numbers of hyperedges. Although this is a hard task, our model consistently outperforms all the baselines significantly, with an average improvement of 28.3% in terms of the F1 score. This demonstrates that our model can learn the hyperedges well under the MIMR objective, without the direct supervision of hyperedge ground truths.
Appendix H More Ablation Studies
H.1 Choices of Number of Hyperedges $K$
@@
Figure 8: Studies on the choice of two key hyperparameters, $\beta$ and $K$, in our model.
@@
H.2 Choices of the Trade-off Coefficient $\beta$
@@
Appendix I Runtime Comparison
@@
Figure 10: Visualization of the most significant hyperedge of the EN-back task.
Appendix J More Visualizations
Resting-State Brain Region Importance
Figure 11: Region importance of resting state.
@@
We visualize the most significant hyperedge of the EN-back task in Figure 10. We observe a coordinated interaction of numerous brain regions, each fulfilling specific roles. Notably, some of these regions serve multi-functional purposes:
@@
  • Visual Processing: Calcarine_R, Lingual_L, Fusiform_L. These regions are responsible for visual perception and some of them are related to complex visual contents like symbols and human faces, which were presented during the fMRI task.